OpenStudy (anonymous):

Find the general term a[n] for the geometric sequence. a[1]=-5 and a[2]=15 a[n] =

hartnn (hartnn):

the general n term formula for geometric sequence is \(a[n] = a[1] r^{n-1} \) where r= common ratio = next term / current term here r= 15/ (-5) = .... ?

OpenStudy (anonymous):

-3

hartnn (hartnn):

yes, so just plug in values of r and a[1] in that formula !

OpenStudy (anonymous):

so then would the correct response be a[n] = -5(-3)^n-1....?

hartnn (hartnn):

yes, that would be correct, just don't miss on brackets a[n] = -5(-3)^{n-1}

hartnn (hartnn):

\(a[n] = -5(-3)^{n-1}\)

OpenStudy (anonymous):

Thank you very much

hartnn (hartnn):

welcome ^_^